OpenStudy (anonymous):

Given the matrices A and B below, find A + B find 3A. A=(2 -3) (0 5) (7 -2) B=(1 0) (10 -1/2) (3 1)

OpenStudy (anonymous):

Are you confused about how to do this?

OpenStudy (anonymous):

yes i dont understand what the question is asking me to do

OpenStudy (anonymous):

For matrix addition: \[\left[\begin{matrix}a & b\\c&d\end{matrix}\right]+\left[\begin{matrix}g & h\\l&k\end{matrix}\right] = \left[\begin{matrix}a+g & b+h\\c+l&d+k\end{matrix}\right]\]

OpenStudy (anonymous):

So just add the corresponding elements and put the sum in the same position as the original elements.

OpenStudy (anonymous):

so then would it be (3 7 ) (10 4.5) (10 -1)

OpenStudy (anonymous):

but in the right format

OpenStudy (anonymous):

Not quite. The first column is correct. The second isnt

OpenStudy (anonymous):

and for the 3A one should i just multiply everything in the A matrace by 3?

OpenStudy (anonymous):

yes, that is right for the multiplication.

OpenStudy (anonymous):

I don't think the first number should be a 7 in column 2 unless you mean for the second matrix to have a 10 not a 0.

OpenStudy (anonymous):

oh right it should be -3

OpenStudy (anonymous):

oh right it should be -3

OpenStudy (anonymous):

yep. That's all there is to it.

OpenStudy (anonymous):

\[A+B=\left( \begin{array}{cc} 3 & -3 \\ 10 & \frac{9}{2} \\ 10 & -1 \end{array} \right)\text{, } 3A=\left( \begin{array}{cc} 6 & -9 \\ 0 & 15 \\ 21 & -6 \end{array} \right) \]